Determine the molar solubility of AgI in pure water.


Ksp (AgI)= 8.51 x 10-17
Answer #1

Ksp=x^2=8.51 x 10^-17

So x=9.22 x 10^-9 M
Answer #2

s^2=8.51*10^-17

So

s=9.225*10^-9 is the molar solubility
Answer #3

let s be the solubility... so s and s dissociates...

=> s^2 = 8.51*10^-17 => s = 9.22*10^-9 M